OpenStudy (anonymous):

find the range r if the quadratic equation x(x-2)=2(x+r) has no real roots

OpenStudy (anonymous):

HINT: Discriminant will be less than 0

OpenStudy (anonymous):

x^2 -2x - 2x -2r = 0 x^2 - 4x - 2r = 0 delta = 4 + 2r < 0 -> r < -2

OpenStudy (anonymous):

x^2 -2x - 2x -2r = 0 x^2 - 4x - 2r = 0 (-4)^2 -4(1) (-2r)<0 16 +8r<0 8r<-16 r<-2.... is it right?

OpenStudy (anonymous):

yes, seems good to my tired eyes.

OpenStudy (anonymous):

haha..Thank You